OpenStudy (anonymous):

Suppose a triangle has two sides of length 3 and 4 and that the angle between these two sides is pi/3. What is the length of the third side of the triangle?

OpenStudy (anonymous):

A. \[4\sqrt{3}\] B. \[\sqrt{3}\] C. 3 D \[\sqrt{13}\]

OpenStudy (anonymous):

Its radial 13